# Conformal e-prediction

### v3

@@ 4 Conditional conformal e-predictors
Conformal prediction and conformal e/prediction satisfy the marginal property of validity that we have discussed so far. In this section we will discuss stronger properties of validity, including what we called object-conditional and label-conditional validity in [28, Sect. 4.6]. Suppose, e.g., we have an algorithm for diagnosing Covid. In object-conditional validity, we divide the objects into separate categories (such as men and women) and require validity inside each of the categories, not just on average. In label-conditional validity, we require validity for each possible label, in this case separate validity for people who have Covid and those who do not. More generally, we can talk about observation-conditional validity, where the categories for which we require separate validity are defined in terms of both objects and labels.
The most standard approach to conditional conformal prediction is Mondrian conformal prediction (see, e.g., [28, Sections 4.6.7–4.6.8]), where the observation space $\mathbf{Z}$ is split into a family, often finite, of disjoint categories, and conformal p-values are computed for each category separately. For a relatively narrow group of conformal predictors, object-conditional validity has been established for overlapping categories by Jung et al. [10]. This has been generalized by Gibbs et al. [6] to what may be called “fuzzy categories”: a category becomes a nonnegative function on the object space $\mathbf{X}$, with categories in the sense of subsets of $\mathbf{X}$ corresponding to the indicator functions on $\mathbf{X}$. In this section we will see that the steps of moving from disjoint to overlapping and then to “fuzzy” categories become very simple and natural in conformal e-prediction.
@@
Let us fix $\kappa\in\mathcal{K}$ and check that (13) is true even conditionally on the $\sigma$-algebra generated by $\kappa(Z_{1}),\dots,\kappa(Z_{n}),\kappa(Z)$, by all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=0$, and by the multiset $B$ consisting of $Z$ and all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=1$. Since, conditionally on $\mathcal{F}$ and inside the event $\kappa(Z)=1$, all orderings of $B$ are equiprobable almost surely, the conditional expectation of $f(Z_{1},\dots,Z_{n},Z)$ is 1 given $\mathcal{F}$ and $\kappa(Z)=1$, which implies (13). ∎
@@
By the strong law of large numbers, the antecedent of (14) holds with probability 0 or 1, depending on whether $\mathbb{P}(\kappa(Z)=1)=0$ or $\mathbb{P}(\kappa(Z)=1)>0$, where $Z$ is any of $Z_{n}$. In the case $\mathbb{P}(\kappa(Z)=1)=0$, (14) holds vacuously, so let us assume $\mathbb{P}(\kappa(Z)=1)>0$. Our goal is to prove that the consequent of (14) holds almost surely.
@@
The proofs of Propositions 6 and 7 show that their assumptions (exchangeability or being IID) can be weakened, similarly to the case of Mondrian conformal prediction [28, Sect. 11.3.6]. On the other hand, the conclusion of Proposition 7 can be strengthened to give an iterated-logarithm result along the lines of Proposition 5.
@@ 5 Split conformal e-predictors
where $F_{z_{1},\dots,z_{n-c}}:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$ is the neural network found from $z_{1},\dots,z_{n-c}$ as training set. (Training might take a long time, but applying to a new object $x$ is typically quick.) And an example of an easily computable normalizing transformation is
@@ 6 Cross-conformal e-predictors
Split conformal e-predictors are often computationally efficient, but their predictive efficiency (to be discussed in detail in the next section) may suffer as compared with “full” conformal predictors discussed in Sections 24, since the latter may be said to use the full training set both as training set proper and as calibration set. The idea behind cross-conformal e-prediction is to combine several split conformal predictors in order to achieve better predictive efficiency.
@@
This follows from Proposition 9 and the arithmetic mean of e-variables being an e-variable (this is obvious and discussed in detail in [32, Sect. 3]). ∎
@@
To compare informally the outputs of cross-conformal predictors [28, Sect. 4.4] and cross-conformal e-predictors, we can use the rough transformation discussed in [32, Remark 2.3]: a p-value of $p$ roughly corresponds to an e-value of $1/p$. Under this transformation, the arithmetic average of e-values corresponds to the harmonic average of p-values, and the harmonic average is always less than or equal to the arithmetic average [8, Theorem 16]. This suggests that cross-conformal e-prediction produces better results than cross-conformal prediction does. In the opposite direction, the arithmetic average of p-values corresponds to the harmonic average of e-values, which again suggests that cross-conformal e-prediction produces better results.
@@
It is easy to combine the ideas of this section and Sect. 4 to design conditional cross-conformal e-predictors (or combine Sections 5 and 4), but we stick to the simplest cases.
@@
Proposition 3 continues to hold for cross-conformal e-predictors and, therefore, it gives its time-wise property of validity in the online mode. However, intuitively, using cross-conformal e-prediction in the online mode defeats the purpose of cross-conformal prediction: once we process $x_{1},y_{1},\dots,x_{n},y_{n}$, we would like to apply the rule that we have found (see item 1 on p. 1) to a large number of new objects rather than just one. There are more complicated settings of “weak teachers” (along the lines of [28, Sect. 3.3]) that combine cross-conformal e-prediction with time-wise validity in useful ways, but we will not discuss them further in this paper.
@@ 7 Predictive efficiency of conformal e-predictors
We developed suitable criteria of efficiency for conformal prediction in [28, Sect. 3.1] (whose notation we will use in this section, except that $e$ will stand for e-values). The idea is that a criterion of efficiency cannot be regarded as suitable if in the limiting case of infinite training and test sets it leads to very unnatural optimal nonconformity measures. An example of such an unsuitable criterion of efficiency for conformal prediction is the use of average confidence and credibility, as defined in the first edition of [28] (analogously to our definition in Sect. 2 above) and analysed in [28, Sections 3.1.6–3.1.7]. This section proposes natural criteria of efficiency for conformal e-prediction that lead to natural nonconformity e-measures.
@@
Let us find the optimal nonconformity e-measure for the observed log criterion in the “idealised setting” akin to the one considered in [28, Sect. 3.1.4] in the case of conformal prediction. For that it will be convenient to modify slightly the definition of a nonconformity e-measure.
@@
As in [28, Sect. 3.1.5], we assume that the object space $\mathbf{X}$ and the label space $\mathbf{Y}$ are finite (the latter simply means that we are interested in pattern recognition). We identify a probability measure $Q$ on a finite set $A$ (such as $\mathbf{Y}$) with a function mapping $a\in A$ to $Q(\{a\})$, often omitting the curly braces in expressions such as $Q(\{a\})$. If $Q$ is a probability measure on $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$, let $Q_{\mathbf{X}}\in\mathbf{P}(\mathbf{X})$ be its marginal probability measure on $\mathbf{X}$, and let $Q_{x}\in\mathbf{P}(\mathbf{Y})$, $x\in\mathbf{X}$, be the conditional probability measure on $\mathbf{Y}$ given $x$:
@@
The rest of this section will be mainly devoted to the proof of Proposition 14. First let us check that (22) is indeed a nonconformity e-measure. It is even even true that its average is 1 over each $P_{x}$:
@@
In the terminology of [28, Sect. 3.1], the criterion of efficiency (19) is “observed” in that the efficiency of the conformal e-predictor on a test observation $z_{i}$ is measured by an expression, namely $\sum_{y\neq y_{i}}\ln e_{i}^{y}$, that depends on the observed true label $y_{i}$. A natural alternative to (19) is the prior log criterion
@@
(We have already used this nonconformity e-measure with $P_{x}$ replaced by its estimate: see (18) above.) Notice that while in conformal prediction suitable (“conditionally proper”) observed and prior criteria of efficiency lead to the same optimal conformity measures [28, Theorem 3.1], in the case of conformal e-prediction the optimal nonconformity e-measures are different, albeit typically very close for a large label space $\mathbf{Y}$.
8 Conformal e-testing
In this and next sections we discuss testing the assumption of exchangeability. We start in this section from batch testing, which is standard in statistics (although we use terminology that is standard in conformal testing rather than statistics). We assume that the number $N$ of observations is fixed and known in advance. At the end of step $N$ we are required to make a decision whether to reject the hypothesis of exchangeability, or at least to present the amount of evidence that we found against the null hypothesis of exchangeability. In the next section we will consider testing the assumption of exchangeability repeatedly, which we will refer to as multistage testing; as soon as we suspect that exchangeability is violated, we raise an alarm, and we are allowed do so more than once.
Conformal e-testing in the batch mode is defined in two steps. First, a function $E:\mathbf{Z}^{N}\to[0,\infty)$ is called a basic conformal e-test if it has the form
$E(z_{1},\dots,z_{N})=\prod_{n=1}^{N}f(z_{1},\dots,z_{n}),\quad\forall(z_{1},\dots,z_{N})\in\mathbf{Z}^{N},$
for some conformal e-predictor $f$. This definition corresponds to the following gambling picture: we start from an initial capital of $1$ and then compound the conformal e-values as usual by multiplying them (ignoring the fact that the $n$th e-value output by $f$ is not an e-value conditionally on $e_{1},\dots,e_{n-1}$). And second, a conformal e-test is defined as a convex combination of basic conformal e-tests: namely, the conformal e-tests are defined as convex mixtures $\lambda_{1}E_{1}+\dots+\lambda_{k}E_{k}$ of basic conformal e-tests $E_{1},\dots,E_{k}$, where $k\in\{1,2,\dots\}$, $\lambda_{i}\in[0,1]$, and $\lambda_{1}+\dots+\lambda_{k}=1$.
Proposition 16.
Suppose $Z_{1},\dots,Z_{N}$ are exchangeable. For any conformal e-test $E$, $E(Z_{1},\dots,Z_{N})$ is a bona fide e-variable. It satisfies $\mathbb{E}(E(Z_{1},\dots,Z_{N}))=1$ if the underlying conformal e-predictor is admissible.
Proof.
Assume, without loss of generality, that $E$ is a basic conformal e-test. Let us set $E_{n}:=f(Z_{1},\dots,Z_{n})$, where $f$ is the underlying conformal e-predictor, and use the exchangeable filtration $(\mathcal{F}_{n})$. We can show that $\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N+1})\leq 1$ a.s. by induction in $N$ (for $N=0$ this statement is vacuously true):
$\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N+1})=\mathbb{E}(\mathbb{E}(E_{1}\dots E_{N}\mid\mathcal{F}_{N})\mid\mathcal{F}_{N+1})\\ =\mathbb{E}(E_{N}\mathbb{E}(E_{1}\dots E_{N-1}\mid\mathcal{F}_{N})\mid\mathcal{F}_{N+1})\leq\mathbb{E}(E_{N}\mid\mathcal{F}_{N+1})\leq 1\quad\text{a.s.}$
If the underlying conformal e-predictor is admissible, both $\leq$” become “$=$”. ∎
Proposition 16 is a statement of validity of conformal e-testing. If $E$ is chosen in advance and takes a very large value on the realized data sequence, we are justified in rejecting the assumption of exchangeability.
As the next step, we explore the efficiency of conformal e-testing along the lines of the treatment of the efficiency of conformal testing in [28, Sect. 9.1]. As in the case of conformal testing [28, Sect. 9.1], in our (rather weak) statement of efficiency we will restrict ourselves to the binary case, $\mathbf{Z}=\{0,1\}$.
First we define unrestricted testing of exchangeability. A probability measure $P$ on $\mathbf{Z}^{N}$ is exchangeable if, for any permutation $\pi$ of $\{1,\dots,N\}$ and any measurable $A\subseteq\mathbf{Z}^{N}$,
$P(A)=P\{(z_{1},\dots,z_{N}):(z_{\pi(1)},\dots,z_{\pi(N)})\in A\}.$
The upper exchangeability probability $\operatorname{\mathbb{P}^{\mathrm{exch}}}$ of measurable $A\subseteq\mathbf{Z}^{N}$ is defined as
$\operatorname{\mathbb{P}^{\mathrm{exch}}}(A):=\sup_{P}P(A),$
$P$ ranging over the exchangeable probability measures on $\mathbf{Z}^{N}$. The intuition behind $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ is that, if it is very small, $A$ can be used for testing the exchangeability of the data-generating distribution: we are entitled to reject exchangeability if $(z_{1},\dots,z_{N})\in A$ provided $A$ is chosen in advance.
The upper conformal e-probability $\operatorname{\mathbb{P}^{\mathrm{ce}}}$ of $A\subseteq\mathbf{Z}^{N}$ is defined as
$\operatorname{\mathbb{P}^{\mathrm{ce}}}(A):=\inf\left\{\alpha:\exists E\;\forall(z_{1},\dots,z_{N})\in A:E(z_{1},\dots,z_{N})\geq 1/\alpha\right\},$
$E$ ranging over the conformal e-tests. The intuition is the same: if $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$ is very small, we are entitled to reject exchangeability if we observe $(z_{1},\dots,z_{N})\in A$, again assuming $A$ is chosen in advance. The difference from $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ is that now the lack of exchangeability must be demonstrated via conformal e-testing.
The following proposition is an analogue of [28, Proposition 9.5]. Intuitively, its second statement says that conformal e-testing is universal in the batch mode: if lack of exchangeability can be demonstrated at all, it can be demonstrated (albeit less convincingly) using conformal e-testing.
Proposition 17.
For any event $A\subseteq\mathbf{Z}^{N}$, $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$. Assuming $\mathbf{Z}=\{0,1\}$, $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq N\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$.
Proof.
The inequality $\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)$ follows immediately from Markov’s inequality applied to conformal e-tests $E$ in combination with Proposition 16: if $E(z_{1},\dots,z_{N})\geq 1/\alpha$ for all $(z_{1},\dots,z_{N})\in A$,
$\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)\leq\mathbb{P}(E(Z_{1},\dots,Z_{N})\geq 1/\alpha)\leq\frac{\mathbb{E}(E(Z_{1},\dots,Z_{N}))}{1/\alpha}\leq\alpha,$
assuming $Z_{1},\dots,Z_{N}$ are exchangeable.
Now assume $\mathbf{Z}=\{0,1\}$; this part of the proof will be a modification of the proof of Proposition 9.5 in [28, Sect. 9.4.2]. As a first step, notice that it suffices to prove $\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq\operatorname{\mathbb{P}^{\mathrm{exch}}}(A)$ for any nonempty $A\subseteq\mathbf{Z}^{N}$ such that each sequence in $A$ has the same number of 1s. Let us fix such an $A$, and let $K\in\{0,\dots,N\}$ be the number of 1s in the elements of $A$. For each sequence $\zeta=(z_{1},\dots,z_{N})\in A$, consider the basic conformal e-test $E_{\zeta}=e_{1}\dots e_{N}$, where the $n$th e-value $e_{n}$ is
$e_{n}:=\begin{cases}n/k&\text{if $z_{n}=1$}\\ n/(n-k)&\text{if $z_{n}=0$},\end{cases}$
$k$ being the number of 1s among the first $n$ elements of $\zeta$. (This was called “reckless gambling” in [28, end of Sect. 9.1.2].) The product $E_{\zeta}=e_{1}\dots e_{N}$ will then be
$\frac{N!}{K!(N-K)!}=\binom{N}{K}.$
The arithmetic mean of $E_{\zeta}$ over $\zeta\in A$ witnesses that
$\operatorname{\mathbb{P}^{\mathrm{ce}}}(A)\leq\lvert A\rvert/\binom{N}{K}=\operatorname{\mathbb{P}^{\mathrm{exch}}}(A).\qed$
9 Multistage exchangeability testing
In the previous section we discussed testing the assumption of exchangeability once, but in some important applications we would like to test it repeatedly over time in an online mode (see, e.g., [28, Sect. 8.3]). As usual, such procedures are required to satisfy properties of validity and efficiency. Validity means that, under exchangeability, the probability or frequency of alarms (which in this case are false alarms) should be bounded by a prespecified constant. Efficiency means that, if exchangeability is violated at some point, an alarm should be raised as quickly as possible afterwards.
In this paper we concentrate on validity of multistage testing; as typical for conformal prediction, we do not have enough structure for efficiency results, unless they are empirical or asymptotic or very weak (as those in the previous section). See [28, Sect. 8.6.3] for a more detailed discussion.
Suppose we observe a sequence of e-values $e_{1},e_{2},\dots$ output by a conformal e-predictor in the online protocol, as described in Sect. 3, under exchangeability. The CUSUM procedure [13] raises the $k$th alarm, $k=1,2,\dots$, at the time
$\tau_{k}:=\min\left\{n>\tau_{k-1}:\max_{i\in\{\tau_{k-1}+1,\dots,n\}}e_{i}\dots e_{n}\geq c\right\},$ (27)
where $\tau_{0}:=0$ and $c>1$ is the parameter of the procedure. It is usually applied in the situation where the observations $z_{i}$ are generated independently first from a known probability measure $Q_{0}\in\mathbf{P}(\mathbf{Z})$ and then from another known probability measure $Q_{1}\in\mathbf{P}(\mathbf{Z})$, and where $e_{i}$ is the likelihood ratio of $Q_{1}$ to $Q_{0}$ evaluated at $z_{i}$ (see, e.g., [16, Sect. 6.2]). In our current context, where the $e_{i}$ are conformal e-values, we may call it the conformal CUSUM e/procedure.
The following proposition gives an asymptotic property of validity for the conformal CUSUM e-procedure.
Proposition 18.
Let $A_{n}$ be the number of alarms
$A_{n}:=\max\{k\mid\tau_{k}\leq n\}$ (28)
raised by the conformal CUSUM e-procedure (27) after processing the first $n$ observations $Z_{1},\dots,Z_{n}$. Then
[…]
provided the observations $Z_{1},Z_{2},\dots$ are exchangeable.
A popular modification of the CUSUM procedure is the Shiryaev–Roberts procedure, which replaces the $\max$ in (27) by the sum $\sum$ (see, e.g., [16, Sect. 6.5]). To state our next validity result (Proposition 19 below), we need the modification
$\tau_{k}:=\min\left\{n>\tau_{k-1}:\max_{i\in\{\tau_{k-1}+1,\dots,n\}}\sum_{j=i}^{n}e_{i}\dots e_{j}\geq c\right\},$ (30)
which we will call the reverse Shiryaev–Roberts procedure (or conformal reverse Shiryaev–Roberts e-procedure to emphasize $e_{i}$ being produced by a conformal e-predictor).
Proposition 19.
Let $A_{n}$ be defined by (28) for the conformal reverse Shiryaev–Roberts e-procedure (30). Then we still have (29) provided the observations $Z_{1},Z_{2},\dots$ are exchangeable.
See Appendix B for the proofs. We will, of course, deduce Proposition 18 from Proposition 19: it is clear that the Shiryaev–Roberts procedure (regular or reverse) raises alarms at least as often as CUSUM does. However, the following example shows an advantage of CUSUM from an intuitive point of view.
Example 20.
Consider the vacuous admissible conformal e-predictor identically equal to 1. The CUSUM procedure based on it will never raise alarms, while the Shiryaev–Roberts e-procedure will raise alarms every $\lceil c\rceil$th step, thereby fully using the permission in the target property of validity (29). This example shows one feature of the Shiryaev–Roberts procedure (shared by the reverse Shiryaev–Roberts procedure) that can be considered its disadvantage: while the CUSUM procedure raises an alarm when it has genuine evidence for disorder, the Shiryaev–Roberts procedure may raise an alarm simply because it is allowed to do so by a given constraint; we might not have any evidence for disorder.

### v4

@@ 3 Conditional conformal e-predictors
Conformal prediction and conformal e/prediction satisfy the marginal property of validity that we have discussed so far. In this section we will discuss stronger properties of validity, including what we called object-conditional and label-conditional validity in [25, Sect. 4.6]. Suppose, e.g., that we have an algorithm for diagnosing Covid. In object-conditional validity, we divide the objects into separate categories (such as men and women) and require validity inside each of the categories, not just on average. In label-conditional validity, we require validity for each possible label, in this case separate validity for people who have Covid and those who do not. More generally, we can talk about observation-conditional validity, where the categories for which we require separate validity are defined in terms of both objects and labels.
The most standard approach to conditional conformal prediction is Mondrian conformal prediction (see, e.g., [25, Sections 4.6.7–4.6.8]), where the observation space $\mathbf{Z}$ is split into a family, often finite, of disjoint categories, and conformal p-values are computed for each category separately. For a relatively narrow group of conformal predictors, object-conditional validity has been established for overlapping categories by Jung et al. [10]. This has been generalized by Gibbs et al. [6] to what may be called “fuzzy categories”: a category becomes a nonnegative function on the object space $\mathbf{X}$, with categories in the sense of subsets of $\mathbf{X}$ corresponding to the indicator functions on $\mathbf{X}$. In this section we will see that the steps of moving from disjoint to overlapping and then to “fuzzy” categories become very simple and natural in conformal e-prediction.
@@
Let us fix $\kappa\in\mathcal{K}$ and check that (13) is true even conditionally on the $\sigma$-algebra $\mathcal{F}$ generated by $\kappa(Z_{1}),\dots,\kappa(Z_{n}),\kappa(Z)$, by all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=0$, and by the multiset $B$ consisting of $Z$ and all $Z_{i}$, $i=1,\dots,n$, with $\kappa(Z_{i})=1$. Since, conditionally on $\mathcal{F}$ and inside the event $\kappa(Z)=1$, all orderings of $B$ are equiprobable (almost surely), the conditional expectation of $f(Z_{1},\dots,Z_{n},Z)$ is 1 given $\mathcal{F}$ and $\kappa(Z)=1$, which implies (13). ∎
@@
By the strong law of large numbers, the antecedent of (14) holds with probability 0 or 1, depending on whether $\mathbb{P}(\kappa(Z)=1)=0$ or $\mathbb{P}(\kappa(Z)=1)>0$, where $Z$ is any of the $Z_{n}$. In the case $\mathbb{P}(\kappa(Z)=1)=0$, (14) holds vacuously, so let us assume $\mathbb{P}(\kappa(Z)=1)>0$. Our goal is to prove that the consequent of (14) holds almost surely.
@@
The proofs of Propositions 6 and 7 show that their assumptions (exchangeability or being IID) can be weakened, similarly to the case of Mondrian conformal prediction [25, Sect. 11.3.6]. On the other hand, the conclusion of Proposition 7 can be strengthened to give an iterated-logarithm result along the lines of Proposition 5.
@@ 4 Split conformal e-predictors
where $F_{z_{1},\dots,z_{n-c}}:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$ is the neural network found from $z_{1},\dots,z_{n-c}$ as training set. (Training might take a long time, but applying the rule to a new object $x$ is typically quick.) And an example of an easily computable normalizing transformation is
@@ 5 Cross-conformal e-predictors
Split conformal e-predictors are often computationally efficient, but their predictive efficiency (to be discussed in detail in the next section) may suffer as compared with “full” conformal e-predictors discussed in Sections 13, since the latter may be said to use the full training set both as training set proper and as calibration set. The idea behind cross-conformal e-prediction is to combine several split conformal predictors in order to achieve better predictive efficiency.
@@
This follows from Proposition 9 and the arithmetic mean of e-variables being an e-variable (this is obvious and discussed in detail in [28, Sect. 3]). ∎
@@
To compare informally the outputs of cross-conformal predictors [25, Sect. 4.4] and cross-conformal e-predictors, we can use the rough transformation discussed in [28, Remark 2.3]: a p-value of $p$ roughly corresponds to an e-value of $1/p$. Under this transformation, the arithmetic average of e-values corresponds to the harmonic average of p-values, and the harmonic average is always less than or equal to the arithmetic average [8, Theorem 16]. This suggests that cross-conformal e-prediction produces better results than cross-conformal prediction does. In the opposite direction, the arithmetic average of p-values corresponds to the harmonic average of e-values, which again suggests that cross-conformal e-prediction produces better results.
@@
It is easy to combine the ideas of this section and Sect. 3 to design conditional cross-conformal e-predictors (or to combine Sections 4 and 3), but we stick to the simplest cases.
@@
Proposition 3 continues to hold for cross-conformal e-predictors and, therefore, it gives its time-wise property of validity in the online mode. However, intuitively, using cross-conformal e-prediction in the online mode defeats the purpose of cross-conformal prediction: once we process $x_{1},y_{1},\dots,x_{n},y_{n}$, we would like to apply the rule that we have found (see item 1 on p. 1) to a large number of new objects rather than just one. There are more complicated settings of “weak teachers” (along the lines of [25, Sect. 3.3]) that combine cross-conformal e-prediction with time-wise validity in useful ways, but we will not discuss them further in this paper.
@@ 6 Predictive efficiency of conformal e-predictors
We developed suitable criteria of efficiency for conformal prediction in [25, Sect. 3.1] (whose notation we will use in this section, except that $e$ will stand for e-values). The idea is that a criterion of efficiency cannot be regarded as suitable if in the limiting case of infinite training and test sets it leads to very unnatural optimal nonconformity measures. An example of such an unsuitable criterion of efficiency for conformal prediction is the use of average confidence and credibility, as defined in the first edition of [25] (analogously to our definition in Sect. 1 above) and analysed in [25, Sections 3.1.6–3.1.7]. This section proposes natural criteria of efficiency for conformal e-prediction that lead to natural nonconformity e-measures.
@@
Let us find the optimal nonconformity e-measure for the observed log criterion in the “idealised setting” akin to the one considered in [25, Sect. 3.1.4] in the case of conformal prediction. For that it will be convenient to modify slightly the definition of a nonconformity e-measure.
@@
As in [25, Sect. 3.1.5], we assume that the object space $\mathbf{X}$ and the label space $\mathbf{Y}$ are finite (the latter simply means that we are interested in pattern recognition). We identify a probability measure $Q$ on a finite set $A$ (such as $\mathbf{Y}$) with a function mapping $a\in A$ to $Q(\{a\})$, thus often omitting the curly braces in expressions such as $Q(\{a\})$. If $Q$ is a probability measure on $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$, let $Q_{\mathbf{X}}\in\mathbf{P}(\mathbf{X})$ be its marginal probability measure on $\mathbf{X}$, and let $Q_{x}\in\mathbf{P}(\mathbf{Y})$, $x\in\mathbf{X}$, be the conditional probability measure on $\mathbf{Y}$ given $x$:
@@
The rest of this section will be mainly devoted to the proof of Proposition 14. First let us check that (22) is indeed a nonconformity e-measure. It is even true that its average is 1 over each $P_{x}$:
@@
In the terminology of [25, Sect. 3.1], the criterion of efficiency (19) is “observed” in that the efficiency of the conformal e-predictor on a test observation $z_{i}$ is measured by an expression, namely $\sum_{y\neq y_{i}}\ln e_{i}^{y}$, that depends on the observed true label $y_{i}$. A natural alternative to (19) is the prior log criterion
@@
(We have already used this nonconformity e-measure with $P_{x}$ replaced by its estimate: see (18) above.) Notice that while in conformal prediction suitable (“conditionally proper”) observed and prior criteria of efficiency lead to the same optimal conformity measures [25, Theorem 3.1], in the case of conformal e-prediction the optimal nonconformity e-measures are different, albeit typically very close for a large label space $\mathbf{Y}$.